OpenStudy (anonymous):

Identify a quadratic function that fits the points (−1, 8),(2, −1), and (3, 0).

OpenStudy (perl):

you can use the standard form of a quadratic y = ax^2 + bx + c Plug in the points given, you will have a system of three equations in three unknowns For example if we plug in (-1,8) we get 8 = a *(-1)^2 +b(-1) + c

OpenStudy (perl):

which simplifies to 8 = a - b + c

OpenStudy (anonymous):

oh okies and what do i do next?

OpenStudy (perl):

plug in the other two points

OpenStudy (perl):

plug (2,-1) into y = ax^2 + bx + c -1 = a(2)^2 + b(2) + c this simplifies to -1 = 4a + 2b + c

OpenStudy (anonymous):

f(x)= x^2-4x+3

OpenStudy (perl):

yes that looks right, how did you get that?

OpenStudy (anonymous):

this online calculator :D

OpenStudy (perl):

do you have a link ?

OpenStudy (perl):

you can also solve it using TI-84